OpenStudy (anonymous):

Ralph is 3 times as old as Sara. In 6 years, Ralph will be only twice as old as Sara will be then. Find Ralph's age now. Which of the following equations could be used to solve the problem? 3x + 2x x + 6 = 2(3x + 6) 3x + 6 = 2(x + 6) 3x + 6 = 2x + 6

OpenStudy (anonymous):

3rd?

OpenStudy (anonymous):

@Callisto

OpenStudy (anonymous):

@confluxepic

OpenStudy (dn_20):

So y = years old of Ralph x = of years old of Sara given: y = 3x using the third function, or C 3x +6 = 2(x +6) 3x = 2x +6 x = 6 Plug in y = 3 (6) y = 18 6 years later: x = 6 +6 = 12 y = 18 + 6 = 24 x+6 = 2x = y+6 So the answer is C

OpenStudy (dn_20):

or the third one, I will say

OpenStudy (anonymous):

thank you:)

OpenStudy (anonymous):

It could also work if sarah was 3 and Ralph was 12

OpenStudy (anonymous):

oops not 3, 4

OpenStudy (anonymous):

NM this entire equation for me is a train wreak :/ sorry lol! just go with the other guys